# Generalized Lotka-Volterra model with sparse interactions: non-Gaussian effects and topological multiple-equilibria phase

### v1

@@ Abstract
We study the equilibrium phases of a generalized Lotka-Volterra model characterized by a species interaction matrix which is random, sparse and symmetric. Dynamical fluctuations are modeled by a demographic noise with amplitude proportional to the effective temperature $T$. The equilibrium distribution of species abundances is obtained by means of the cavity method and the Belief Propagation equations, which allow for an exact solution on sparse networks. Our results reveal a rich and non-trivial phenomenology that deviates significantly from the predictions of fully connected models. Consistently with data from real ecosystems, which are characterized by sparse rather than dense interaction networks, we find strong deviations from Gaussianity in the distribution of abundances. In addition to the study of these deviations from Gaussianity, which are not related to multiple-equilibria, we also identified a novel topological glass phase, present at both finite temperature, as shown here, and at $T=0$, as previously suggested in the literature. The peculiarity of this phase, which differs from the multiple-equilibria phase of fully-connected networks, is its strong dependence on the presence of extinctions. These findings provide new insights into how network topology and disorder influence ecological networks, particularly emphasizing that sparsity is a crucial feature for accurately modeling real-world ecological phenomena.
@@ I Introduction
Since the pioneering work of May [1], recent advances in statistical mechanics have opened new avenues for understanding the stability and phase behaviour of large ecosystems. In particular, the generalized Lotka-Volterra (gLV) model has emerged as a powerful framework for studying the coexistence of species in complex ecological networks. In the presence of symmetric and dense interactions it is possible to exactly solve the equilibrium statistical mechanics, characterizing completely the phase diagram of the system. In general the equilibrium phases of the gLV model with disordered interactions are controlled by three parameters: $\mu$, the average interaction strength; $\sigma$, the width of the disordered interactions distribution; $T$, the amplitude of thermal fluctuations. Typically, in dense networks, at finite values of $\mu$ and $T$, a multiple-equilibria glass phase emerges for large values of $\sigma$ [2]. However, realistic ecological networks are neither dense nor characterized by symmetric interactions: it is therefore legitimate to wonder which aspects of the phenomenology of these exactly solvable models represent features of realistic ecosystems and which ones are simply due to model approximations. While several papers already considered the case of asymmetric interactions [3, 4, 5, 6, 7], here we focus on symmetric ones, so that we can exploit the tools of equilibrium statistical mechanics. We consider interactions which have quenched disorder, the case of annealed one being discussed elsewhere [8], and which are sparse, at variance with most of the previous literature. In particular we will focus on the study of species abundance distributions, which in sparse networks may have very relevant fluctuations. It is in fact well known that species abundance distributions found from real ecosystems data are typically log-normal or gamma-like distributed [9, 10, 11, 12, 13, 14]; however, in the case of normally distributed quenched interactions, truncated Gaussians are the standard for dense networks [3, 4, 15]. Whilst deviations from Gaussianity can be obtained by assuming non-Gaussian quenched disorder [16], the goal of this work is to show that non-Gaussian effects can also emerge owing to sparsity, while keeping Gaussian quenched disorder.
@@
The main focus of our analysis will be to understand the behaviour of the gLV model on a Random Regular Graph, which we will also refer to as Bethe Lattice. Specifically, we consider graphs with small connectivity $k$, meaning that each species interacts with $k$ other species selected at random from the $N$ possible ones (in the following, we will also refer to the connectivity of a species as its degree). We will study the properties of the gLV model on a sparse network varying three parameters: the mean value $\mu$ and the standard deviation $\sigma$ of the nonzero normally distributed symmetric interactions and the demographic noise strength $T$. We find two main results: strong deviation from Gaussianity in abundance distributions already in the single equilibrium phase and, for large enough values of the parameter $\mu$, a topological glass phase at finite temperature. This phase looks very different from the glass phase of fully connected models; in particular, as first noticed in [17], in this phase glassiness is strongly correlated to the presence of extinctions. The gLV model on a random regular graph was in fact already considered in [17], where a phase transition to multiple equilibria was studied at zero temperature, $T=0$, zero disorder, $\sigma=0$, and finite interaction strength $\mu$. Here we extend a similar analysis to finite temperatures.
@@
The work will be organized as follows: In Section II we will introduce the model and present the so called cavity equations used to obtain the equilibrium abundance marginal distributions; in Section III we will discuss in detail the deviations from Gaussianity which can be found in the abundance distributions on a sparse ecological network. We will then show that no glass phase emerges at small values of $\mu$, at variance with fully-connected models. Sec. IV will be then devoted to study the equilibrium phase diagram of the model at zero disorder ($\sigma=0$), varying only $\mu$ and $T$. The equilibrium properties in this regime have been investigated using both the cavity equations and the Langevin dynamics, finding consistency between the two methods. Conclusions and perspectives will then be reported in Sec. V.
@@ II Generalized Lotka-Volterra model on the Bethe lattice
where $n_{i}(t)$ is the abundance of species ($i=1,...,N$) at time $t$ and $\partial i$ represents the set of species which interact with species $i$. We will focus on random graphs for which the cardinality of $\partial i$ is fixed and equal to $k_{i}=3$ for all $i$’s, even if the methods we use could be applied with no modification also in the case of a poissonian distribution of the connectivities $k_{i}$’s. The factors $r_{i}$ and $K_{i}$ are respectively the intrinsic growth rate and the carrying capacity of species $i$, while $\xi_{i}(t)$ represents a Gaussian multiplicative noise with zero mean and covariance
@@
We consider a symmetric interaction matrix with elements $\alpha_{ij}=\alpha_{ji}$. Because of the symmetric interactions we know that the Langevin dynamics admits an equilibrium distribution of the form $P({\bf n})=\exp(-H_{\text{eff}}({\bf n})/T)$, with ${\bf n}=\{n_{1},\ldots,n_{N}\}$ [2, 20]. When the stochastic dynamics of Eq. (1) is complemented with a reflecting wall condition $n_{i}^{\text{min}}=\lambda$ with $\lambda\ll 1$ for every species $i$, used to avoid unphysical negative values of abundances in numerical simulations, the effective Hamiltonian $H_{\text{eff}}({\bf n})$ reads as:
@@
where the term with the Heaviside function $\theta({n}_{i}-\lambda)$ accounts for the presence of the reflecting wall. The symbol $E$ in the double sum of Eq. (II) denotes the set of pairs of interacting species, which in the present work are arranged in a sparse random graph. Without any lack of generality, we will consider identical carrying capacities, $K_{i}=K$, and identical intrinsic growth rates, $r_{i}=r$, for all species. In particular we choose $r=1$ and $K=280$, where the choice of a large value of $K$ allows to better appreciate the shape of abundance distributions. We study a sparse network with finite connectivity $k=3$, where the variables $n_{i}$ take only discrete non-negative values, consistently with the interpretation of $n_{i}$ as the number of individuals for the species $i$. In this particular case, the choice of discrete variables is also an algorithmic necessity for an efficient computational way to solve the Belief Propagation equations, which we will present in the following. Taking into account the discreteness of $n_{i}$ values, we modify the effective Hamiltonian of Eq. (II): the term $\ln\theta(n_{i}-\lambda)$ is dropped and the term $T\ln(n_{i})$ is replaced by $T\ln(n_{i}+\epsilon)$, with $\epsilon$ small but finite ($\epsilon=0.0001$) in order to regularize the distribution at $n_{i}=0$. The elements of the symmetric matrix $\alpha_{ij}$ are taken from a Gaussian distribution with mean and variance which, consistently with the literature on disordered systems [2, 19], are defined as:
@@
In the following we will use the so-called cavity method [21] to sample the equilibrium Boltzmann distribution $P({\bf n})\propto e^{-\beta H_{\alpha}({\bf n})}$ of the discrete variables $n_{i}$ and in particular to compute their marginals. The crucial observation is that a random regular graph is a locally tree-like graph, for which the typical length of a loop is $\log(N)$ and thus it diverges in the large-$N$ limit (as also for a graph with a Poissonian degree distribution). The main property of a random regular graph is therefore that in the thermodynamic limit, due to the absence of loops, local marginals become effectively factorized. It is in this perspective that, given two neighbouring nodes $i$ and $j$, one introduces the so called cavity marginal $\eta^{\alpha}_{i\rightarrow j}(n_{i})$, where $\alpha$ denotes an instance of the quenched disorder. The cavity marginal is defined as the marginal probability distribution of the variable $n_{i}$ in a graph where the edge connecting the node $i$ with the node $j$ has been removed. The cavity marginals will satisfy the following self-consistent equations, that are called Belief-Propagation (BP) equations:
@@
and where we indicate with $\partial i\backslash j$ the set of neighbours of $i$, excluding node $j$. The solutions of Eqs. (II), (II) can be obtained iteratively: details on the derivation of the equations and the implementation of the algorithmic procedure are given in Appendix A. In particular we have run the BP algorithm on a given graph, where each graph, being the connectivity fixed, is completely characterized by the choice of the quenched disordered couplings. At convergence we have a number $2\absolutevalue{E}$ | of cavity marginals $\eta^{\alpha}_{i\rightarrow j}$, where |$\absolutevalue{E}$ | is the number of non-oriented edges. Once the cavity marginals have been computed, one can extract from them the marginal distribution $\eta^{\alpha}_{i}(n_{i})$ for each species $i$ as:
@@
where $z^{\alpha}_{i}$ is a normalization factor ensuring that $\sum_{n_{i}}\eta^{\alpha}_{i}(n_{i})=1$. Note that, because of the quenched disorder and the sparsity of the interaction network, the marginals for different species will be different. This is a crucial difference with respect to the fully-connected case where, for a given instance of the disorder, all marginals are identical by construction. Taking into account that on sparse graphs the marginals may have fluctuations not only with respect to the graph realization but also from species to species, we consider two kinds of average for the distribution $\eta^{\alpha}_{i}(n_{i})$: the sample average,
@@
III Non-Gaussianity in the single-equilibrium phase
The disordered system approach to large ecosystems has provided many interesting and suggestive insights on their behaviour and their response/stability with respect to external perturbations [1, 22, 2, 23]. Nevertheless, the robustness of the exact results obtained so far for the gLV problem depends on a very crude approximation in the model: the consideration of fully-connected interaction networks between species [3, 4], something which is clearly very unrealistic from the ecological systems perspective. A few points need to be carefully investigated: which aspects of the phenomenology obtained so far by exact results from disordered systems techniques are consequences only of the fully-connected networks? Are there features intrinsically related to the sparsity of ecological networks, as for instance fat-tailed or otherwise non-Gaussian tailed abundance distributions, which cannot be reproduced by the fully-connected model with normally distributed random interactions? In addition to addressing these questions, our work aims to determine whether the glass phases found in fully connected models remain robust in sparse networks, making them more likely to occur in realistic ecosystems.
@@ III Non-Gaussianity in the single-equilibrium phase
Figure 1: Species abundance average marginal distribution $\eta(n)$ for $T=1$ and $\hat{\mu}=0.1$, for $\hat{\sigma}=0.02$ (blue) and $\hat{\sigma}=0.20$ (orange) (Both in the single equilibrium phase). Solid lines are fits for the data: the black line is a Gaussian fit for $\hat{\sigma}=0.02$, the red line is a fit with a Gamma distribution as in (12), with parameters $\alpha=7.5$ and $\beta=0.029$ plus a gaussian peaked at $n=0$ for $\hat{\sigma}=0.20$. Inset: marginal at $\hat{\sigma}=0.20$ computed using BP (orange points) and from the dynamics (green points). The two agree very well. The number of species considered is $N=256$.
In Fig. 1 we show the averaged species abundance marginal distribution $\eta(n)$, Eq.(11), at $T=1$ and $\hat{\mu}=0.1$, for two different values of the standard deviation of the disordered couplings: $\hat{\sigma}=0.02$ and $\hat{\sigma}=0.20$. By studying the convergence of the Belief Propagation algorithm, which will be discussed more thoroughly in Sec. III.1, we know that both these two values of the parameter $\hat{\sigma}$ correspond to a single equilibrium phase. The remarkable finding is that, despite the lack of any glassy feature, upon increasing $\hat{\sigma}$ we find a crossover from a low disorder phase, where the mean marginal $\eta(n)$ is Gaussian, to a high disorder phase where $\eta(n)$ is highly non-Gaussian and is well described by a Gamma distribution. In particular, for $\hat{\sigma}=0.20$, the distribution $\eta(n)$ is well fitted by the red curve in Fig. 1, given by:
@@
where $\Gamma(\alpha)$ is the Gamma function and the fitted parameters are $\alpha=7.5$ and $\beta=0.029$. This is a very significant result as real ecosystems data exhibit species abundance distributions that follow a Gamma distribution [13, 11]. As we show in the top-right inset of Fig. 1, the shape of the non-Gaussian distribution is accurately reproduced by running the Langevin dynamics. This consists in fixing a disordered graph, running the set of equations given by Eq. 1 and stopping the simulation at a sufficiently large time $t_{\text{max}}$ to collect all the $n_{i}(t_{\text{max}})$ for every $i$. By repeating this process for multiple disorder realizations and combining all $n_{i}(t_{\text{max}})$ from each realization, we construct the normalized histogram of species abundances, shown in green in the inset of Fig. 1. Further details on the implementation of the dynamics are provided in Sec. IV.2.
Figure 2: Kurtosis $\kappa(\hat{\sigma},T)$ as a function of standard deviation $\hat{\sigma}$ and temperature $T$ for fixed $\hat{\mu}=0.1$ and $N=256$. For high $\hat{\sigma}$ it is evident that $\kappa>0$ implying that the marginals distributions are developing a non-Gaussian tail.
The crossover from Gaussian to non-Gaussian behaviour varying both the disorder strength $\hat{\sigma}$ and the temperature $T$ is represented in the phase diagram shown in Fig. 2, where color code represents the kurtosis $\kappa(\hat{\sigma},T)$ of the marginal distribution
$\kappa(\hat{\sigma},T)\equiv\frac{<(n-<n>)^{4}>}{<(n-<n>)^{2}>^{2}}-3,$ (13)
where the averages <>$\cdot$>< ⋅ > are taken with respect to $\eta(n)$. For high values of $\hat{\sigma}$ it can be easily recognized the strong deviation from Gaussianity due to the non-gaussian tail of $\eta(n)$ shown in Fig. 1.
@@
The transition from Gaussian to non-Gaussian behaviour is quite robust: it does not only affect the single variable marginal distribution but has also an influence on correlations. This fact is evident when studying the joint probability distribution $P^{\alpha}_{ij}(n_{i},n_{j})$ of a pair of nearest neighbour species on a given graph, which is defined as:
@@
The contour plots showing the behaviour of $P(n,m)$ in the plane $(n,m)$ are reported on the two left panels of Fig. 3 for two different values of the standard deviation: $\hat{\sigma}=0.02$, in the Gaussian regime, and $\hat{\sigma}=0.20$, in the non-Gaussian regime. On the right panels of Fig. 3 we show the contour plot of $P(n,m)-\eta(n)\eta(m)$ for the same two different values of $\hat{\sigma}$. It is clear that $P(n,m)\neq\eta(n)\eta(m)$ and the presence of either correlations or anticorrelations can be fully appreciated. This is an important difference with respect to the fully-connected case where there are no correlations between different species.
Figure 4: Normalized sections of $P(n,n)$ and $\eta(n)\eta(n)$ for $\hat{\sigma}=0.02$, Gaussian regime, and $\hat{\sigma}=0.20$, non-Gaussian regime. We have non trivial correlation effects differently from the fully connected case for which $P(n,n)=\eta(n)\eta(n)$. Also here $\hat{\mu}=0.1$, $T=1$ and $N=256$.
In Fig. 4 we compare $P(n,m)$ for $n=m$ with $\eta(n)\eta(n)$ in order to highlight again the relevance of correlations. In particular, looking at the case $\hat{\sigma}=0.20$ in Fig. 4, we see that $P(n,n)<\eta(n)\eta(n)$ at small values of $n$, which means that positive (competitive) interactions dominate the average correlation. On the contrary we find that for large values of the abundances one has $P(n,n)>\eta(n)\eta(n)$: in this regime negative (mutualistic) interactions are the dominant ones. From the study of the connected probability $P(n,n)-\eta(n)\eta(n)$ it is therefore clear that non-Gaussianity affects (in a non trivial way) also the correlations when the disorder parameter $\hat{\sigma}$ is increased. We have then studied the function $P(n,m)-\eta(n)\eta(m)$ in the specific cases of either exclusively positive ($\alpha_{ij}>0$) or exclusively negative ($\alpha_{ij}<0$) interactions. Namely we have specialized the average of $P^{\alpha}(n,m)$ with respect to the sign of the interaction:

### v2

@@ Abstract
We study the equilibrium phases of a generalized Lotka-Volterra model characterized by a species interaction matrix which is random, sparse and symmetric. Dynamical fluctuations are modeled by a demographic noise with amplitude proportional to the effective temperature $T$. The equilibrium distribution of species abundances is obtained by means of the cavity method and the Belief Propagation equations, which allow for an exact solution on sparse networks. Our results reveal a rich and non-trivial phenomenology that deviates significantly from the predictions of fully connected models. Consistently with data from real ecosystems, which are characterized by sparse rather than dense interaction networks, we find strong deviations from Gaussianity in the distribution of abundances. In addition to the study of these deviations from Gaussianity, which are not related to multiple-equilibria, we also identified a novel topological multiple-attractor phase, present at both finite temperature, as shown here, and at $T=0$, as previously suggested in the literature. The peculiarity of this phase, which differs from the multiple-equilibria phase of fully-connected networks, is its strong dependence on the presence of extinctions. These findings provide new insights into how network topology and disorder influence ecological networks, particularly emphasizing that sparsity is a crucial feature for accurately modeling real-world ecological phenomena.
@@ I Introduction
Since the pioneering work of May [17], recent advances in statistical mechanics have opened new avenues for understanding the stability and phase behaviour of large ecosystems. In particular, the generalized Lotka-Volterra (gLV) model has emerged as a powerful framework for studying the coexistence of species in complex ecological networks. In the presence of symmetric and dense interactions it is possible to exactly solve the equilibrium statistical mechanics, characterizing completely the phase diagram of the system. In general the equilibrium phases of the gLV model with disordered interactions are controlled by three parameters: $\mu$, the average interaction strength; $\sigma$, the width of the disordered interactions distribution; $T$, the amplitude of demographic fluctuations. Typically, in dense networks, at finite values of $\mu$ and $T$, a multiple-equilibria phase emerges for large values of $\sigma$ [1]. However, realistic ecological networks are neither dense nor characterized by symmetric interactions. While the effort to optimize ecosystems modeling is still in progress [5, 26], it is legitimate to wonder which aspects of the phenomenology of fully-connected networks represent features of realistic ecosystems and which ones are simply due to model approximations. While several papers already considered the case of asymmetric interactions [7, 10, 24, 25, 15], here we focus on symmetric ones, so that we can exploit the tools of equilibrium statistical mechanics. We consider interactions which have quenched disorder, the case of annealed one being discussed elsewhere [28], and which are sparse, at variance with most of the previous literature. In particular we will focus on the study of species abundance distributions, which in sparse networks may have very relevant fluctuations. It is in fact well known that species abundance distributions found from real ecosystems data are typically log-normal or gamma-like distributed [18, 27, 30, 4, 13, 19]; however, in the case of normally distributed quenched interactions, truncated Gaussians are the standard for dense networks [7, 10, 11]. Whilst deviations from Gaussianity in the marginal distributions of the single degrees of freedom can be obtained by assuming non-Gaussian quenched disorder [3], and in particular long-tail distributions for the couplings [21], the goal of this work is to show that strong non-Gaussian effects can also emerge owing to sparsity, while keeping Gaussian quenched disorder.
@@
The main focus of our analysis will be to understand the behaviour of the gLV model on a Random Regular Graph, which we will also refer to as Bethe Lattice. Specifically, we consider graphs with small connectivity $k$, meaning that each species interacts with $k$ other species selected at random from the $N$ possible ones (in the following, we will also refer to the connectivity of a species as its degree). We will study the properties of the gLV model on a sparse network varying three parameters: the mean value $\mu$ and the standard deviation $\sigma$ of the nonzero normally distributed symmetric interactions and the demographic noise strength $T$. We find two main results: strong deviation from Gaussianity in abundance distributions already in the single equilibrium phase and, for large enough values of the parameter $\mu$, a topological multiple attractor phase at finite temperature and $\sigma=0$. This phase looks very different from the one of fully connected models; in particular, as first noticed in [16], in this phase the multiple equilibria are linked to different interaction network topologies of surviving species. The gLV model on a random regular graph was in fact already considered in [16], where a phase transition to multiple equilibria was studied at zero temperature, $T=0$, zero disorder, $\sigma=0$, and finite interaction strength $\mu$. Here we extend a similar analysis to finite temperatures.
@@
The work will be organized as follows: In Section II we will introduce the model and present the so called cavity equations used to obtain the equilibrium abundance marginal distributions; in Section III we will discuss in detail the strong deviations from Gaussianity which can be found in the abundance distributions on a sparse ecological network. We will then show that no multiple equilibria phase emerges at small values of $\mu$, at variance with fully-connected models. Sec. IV will be then devoted to study the equilibrium phase diagram of the model at zero disorder ($\sigma=0$), varying only $\mu$ and $T$. The equilibrium properties in this regime have been investigated using both the cavity equations and the Langevin dynamics, finding consistency between the two methods. Conclusions and perspectives will then be reported in Sec. V.
@@ II Generalized Lotka-Volterra model on the Bethe lattice
where $n_{i}(t)$ is the abundance of species (i=1,,Ni=1,...,N) at time $t$ and $\partial i$ represents the set of species which interact with species $i$. We will focus on random graphs for which the cardinality of $\partial i$ is fixed and equal to $k_{i}=3$ for all $i$’s, even if the methods we use could be applied with no modification also in the case of a poissonian distribution of the connectivities $k_{i}$’s. The factors $r_{i}$ and $K_{i}$ are respectively the intrinsic growth rate and the carrying capacity of species $i$, while $\xi_{i}(t)$ represents a Gaussian multiplicative noise with zero mean and covariance
@@
We consider a symmetric interaction matrix with elements $\alpha_{ij}=\alpha_{ji}$. Because of the symmetric interactions we know that the Langevin dynamics admits an equilibrium distribution of the form $P({\bf n})=\exp(-H_{\text{eff}}({\bf n})/T)$, with ${\bf n}=\{n_{1},\ldots,n_{N}\}$ where the Hamiltonian $H_{\text{eff}}({\bf n})$ has been firstly derived in ref. [6] following the Itô convention. When the stochastic dynamics of Eq. (1) is complemented with a reflecting wall condition $n_{i}^{\text{min}}=\lambda$ with $\lambda\ll 1$ for every species $i$, used to avoid unphysical negative values of abundances in numerical simulations, the effective Hamiltonian $H_{\text{eff}}({\bf n})$ reads as:
@@
where the term with the Heaviside function $\theta({n}_{i}-\lambda)$ accounts for the presence of the reflecting wall. The symbol $E$ in the double sum of Eq. (II) denotes the set of pairs of interacting species, which in the present work are arranged in a sparse random graph. Without any lack of generality, we will consider identical carrying capacities, $K_{i}=K$, and identical intrinsic growth rates, $r_{i}=r$, for all species. In particular we choose $r=1$ and $K=280$. We study a sparse network with finite connectivity $k=3$, where the variables $n_{i}$ take only integer non-negative values, consistently with the interpretation of $n_{i}$ as the number of individuals for the species $i$. In this particular case, the choice of discrete variables is also an algorithmic necessity for an efficient computational way to solve the Belief Propagation equations, which we will present in the following and in detail in Appendix A. Please note, however, that the choice to take integer values for $n_{i}$ is not inevitable: indeed one can choose to discretize $n_{i}$ into small steps with a discretization step $dn<1$. We do not expect that the integer discretization influences the results because the discretization step, that in our case is $dn=1$, should be compared with the natural scale for $n$ that is $K$ (fixed point of the single species dynamics) that we chose to be large, $K=280$, exactly for this reason. The independence of the results on the discretization choice is shown in Appendix B. Taking into account the discreteness of $n_{i}$ values, we modify the effective Hamiltonian of Eq. (II): the term $\ln\theta(n_{i}-\lambda)$ is dropped and the term $T\ln(n_{i})$ is replaced by $T\ln(n_{i}+\epsilon)$, with $\epsilon$ small but finite ($\epsilon=0.0001$) in order to regularize the distribution at $n_{i}=0$. In Appendix B we show that the obtained species abundances are almost independent on the exact value of $\epsilon$ as long as it is small enough. The elements of the symmetric matrix $\alpha_{ij}$ are taken from a Gaussian distribution with mean and variance which, consistently with the literature on disordered systems [1, 23], are defined as:
@@
In the following we will use the so-called cavity method [20] to sample the equilibrium Boltzmann distribution $P({\bf n})\propto e^{-\beta H_{\alpha}({\bf n})}$ of the discrete variables $n_{i}$ and in particular to compute their marginals. The crucial observation is that a random regular graph is a locally tree-like graph, for which the typical length of a loop is $\log(N)$ and thus it diverges in the large-$N$ limit (as also for a graph with a Poissonian degree distribution). The main property of a random regular graph is therefore that in the thermodynamic limit, due to the absence of loops, local marginals become effectively factorized. It is in this perspective that, given two neighbouring nodes $i$ and $j$, one introduces the so called cavity marginal $\eta^{\alpha}_{i\rightarrow j}(n_{i})$, where $\alpha$ denotes an instance of the quenched disorder. The cavity marginal is defined as the marginal probability distribution of the variable $n_{i}$ in a graph where the edge connecting the node $i$ with the node $j$ has been removed. The cavity marginals satisfy the following self-consistent equations, that are called Belief-Propagation (BP) equations:
@@
and where we indicate with $\j\partial i\backslash$ j the set of neighbours of $i$, excluding node $j$. The solutions of Eqs. (II), (II) can be obtained iteratively: details on the derivation of the equations and the implementation of the algorithmic procedure are given in Appendix A. In particular we have run the BP algorithm on a given graph, where each graph, being the connectivity fixed, is completely characterized by the choice of the quenched disordered couplings. At convergence we have a number $|2\absolutevalue{E}$ of cavity marginals $\eta^{\alpha}_{i\rightarrow j}$, where $|\absolutevalue{E}$ is the number of non-oriented edges. Once the cavity marginals have been computed, one can extract from them the marginal distribution $\eta^{\alpha}_{i}(n_{i})$ for each species $i$ as:
@@
where $z^{\alpha}_{i}$ is a normalization factor ensuring that $\sum_{n_{i}}\eta^{\alpha}_{i}(n_{i})=1$. These are the exact marginal distributions associated to the equilibrium Gibbs-Boltzmann probability for a locally tree-like graph (as a random-regular graph or a Poissonian graph) as long as the whole Gibbs measure is composed of just one state. Note that, because of the quenched disorder and the sparsity of the interaction network, the marginals for different species will be different. This is a crucial difference with respect to the fully-connected case where, for a given instance of the disorder, all marginals are identical by construction. Taking into account that on sparse graphs the marginals may have fluctuations not only with respect to the graph realization but also from species to species, we consider two kinds of average for the distribution $\eta^{\alpha}_{i}(n_{i})$: the sample average,
@@
III Strong Non-Gaussianity in the single-equilibrium phase
The disordered system approach to large ecosystems has provided many interesting and suggestive insights on their behaviour and their response/stability with respect to external perturbations [17, 9, 1, 22]. Nevertheless, the robustness of the exact results obtained so far for the gLV problem depends on a very crude approximation in the model: the consideration of fully-connected interaction networks between species [7, 10], something which is clearly very unrealistic from the ecological systems perspective. A few points need to be carefully investigated: which aspects of the phenomenology obtained so far by exact results from disordered systems techniques are consequences only of the fully-connected networks? Are there features intrinsically related to the sparsity of ecological networks, as for instance fat-tailed or otherwise non-Gaussian tailed abundance distributions, which cannot be reproduced by the fully-connected model with normally distributed random interactions? In addition to addressing these questions, our work aims to determine whether the multiple attractor phases found in fully connected models remain robust in sparse networks, making them more likely to occur in realistic ecosystems.
@@ III Strong Non-Gaussianity in the single-equilibrium phase
Figure 1: Species abundance average distribution for two values of the coupling variance $\hat{\sigma}$ in the single equilibrium phase. For small $\hat{\sigma}=0.02$ (blue points) the distribution follows a Gaussian distribution (black line), while for large $\hat{\sigma}=0.20$ (orange points) the species abundance follows a Gamma distribution (red line) as in (12), with parameters $\alpha=7.5$ and $\beta=0.029$ plus a gaussian peaked at $n=0$. Inset: marginal at $\hat{\sigma}=0.20$ computed using BP (orange points) and from the dynamics (green points). The two agree very well. In both plots the parameters are $T=1$, $\hat{\mu}=0.1$, $N=256$.
In Fig. 1 we show the averaged species abundance marginal distribution $\eta(n)$, Eq. (11), at $T=1$ and $\hat{\mu}=0.1$, for two different values of the standard deviation of the disordered couplings: $\hat{\sigma}=0.02$ and $\hat{\sigma}=0.20$. By studying the convergence of the Belief Propagation algorithm, which will be discussed more thoroughly in Sec. III.1, we know that both these two values of the parameter $\hat{\sigma}$ correspond to a single equilibrium phase. The remarkable finding is that, despite the lack of any multiple equilibria, upon increasing $\hat{\sigma}$ we find a crossover from a low disorder phase, where the mean marginal $\eta(n)$ is well described by a Gaussian, to a high disorder phase where $\eta(n)$ is highly non-Gaussian and is well described by a Gamma distribution. In particular, for $\hat{\sigma}=0.20$, the distribution $\eta(n)$ is well fitted by the red curve in Fig. 1, given by:
@@
where $\Gamma(\alpha)$ is the Gamma function and the fitted parameters are $\alpha=7.5$ and $\beta=0.029$. This is a very significant result as real ecosystems data exhibit species abundance distributions that follow a Gamma distribution [13, 30].
As we show in the top-right inset of Fig. 1, the shape of the fat-tailed distribution is accurately reproduced by running the Langevin dynamics. This consists in fixing a disordered graph, running the set of equations given by Eq. 1 and stopping the simulation at a sufficiently large time $t_{\text{max}}$ to collect all the $n_{i}(t_{\text{max}})$ for every $i$. By repeating this process for multiple disorder realizations and combining all $n_{i}(t_{\text{max}})$ from each realization, we construct the normalized histogram of species abundances, shown in green in the inset of Fig. 1. Further details on the implementation of the dynamics are provided in Sec. IV.2, while in Appendix D we show that the marginals extracted from BP coincide with those extracted from the dynamics in all the region of parameters for which there is a unique fixed point.
Moreover, we show in Appendix E that in the region of the parameters where the marginals exhibit strong non-gaussianity, variances of the single species abundance distributions follow a power-law dependence on the means of the distributions, well known as Taylor’s law, documented extensively in real ecosystems (see [12]). In this region of parameters, what predicted by our sparse LV model is thus in perfect agreement with what observed in real ecosystems.
Figure 2: Kurtosis $\kappa(\hat{\sigma},T)$ of the average species abundance in the unique fixed point phase. For high variance of the couplings $\hat{\sigma}$ it is evident that $\kappa>0$ implying that the marginals distributions are developing a non-Gaussian tail, while the kurtosis is almost independent on the temperature $T$. The parameters of the model are $\hat{\mu}=0.1$ and $N=256$.
The crossover from almost Gaussian to strongly non-Gaussian behaviour varying both the disorder strength $\hat{\sigma}$ and the temperature $T$ is represented in the phase diagram shown in Fig. 2, where color code represents the kurtosis $\kappa(\hat{\sigma},T)$ of the marginal distribution
$\kappa(\hat{\sigma},T)\equiv\frac{\langle(n-\langle n\rangle)^{4}\rangle}{\langle(n-\langle n\rangle)^{2}\rangle^{2}}-3,$ (13)
where the averages $\langle\cdot\rangle$ are taken with respect to $\eta(n)$. For high values of $\hat{\sigma}$ it can be easily recognized the strong deviation from Gaussianity due to the non-gaussian tail of $\eta(n)$ shown in Fig. 1. Looking at the marginal distribution form in eq. (II), it is evident that there is a factor $\frac{1}{n_{i}+\epsilon}$, associated to the demographic noise, preventing $\eta_{i}^{\alpha}$ from being a perfect Gaussian even in the absence of all the interactions; thus one could be tempted to claim that this could be the reason for the non-Gaussian beaviour of the marginals. In Appendix F we show that indeed the strong deviation from Gaussianity is mainly due to the sparse interactions and not to the demographic noise factor.
@@
The transition from almost Gaussian to strongly non-Gaussian behaviour does not only affect the single variable marginal distribution but has also an influence on correlations. This fact is evident when studying the joint probability distribution $P^{\alpha}_{ij}(n_{i},n_{j})$ of a pair of nearest neighbour species on a given graph, which is defined as:
@@
The contour plots showing the behaviour of $P(n,m)$ in the plane $(n,m)$ are reported on the two left panels of Fig. 3 for two different values of the standard deviation: $\hat{\sigma}=0.02$, in the Gaussian regime, and $\hat{\sigma}=0.20$, in the strongly non-Gaussian regime. On the right panels of Fig. 3 we show the contour plot of $P(n,m)-\eta(n)\eta(m)$ for the same two different values of $\hat{\sigma}$. It is clear that $P(n,m)\neq\eta(n)\eta(m)$ and the presence of either correlations or anticorrelations can be fully appreciated. This is an important difference with respect to the fully-connected case where there are no correlations between different species.
Figure 4: Normalized sections of $P(n,n)$ and $\eta(n)\eta(n)$ for $\hat{\sigma}=0.02$, Gaussian regime, and $\hat{\sigma}=0.20$, non-Gaussian regime. Non trivial correlation effects are evident, differently from the fully connected case for which $P(n,n)=\eta(n)\eta(n)$. The parameters of the model are $\hat{\mu}=0.1$, $T=1$ and $N=256$.
In Fig. 4 we compare $P(n,m)$ for $n=m$ with $\eta(n)\eta(n)$ in order to highlight again the relevance of correlations. In particular, looking at the case $\hat{\sigma}=0.20$ in Fig. 4, we see that $P(n,n)<\eta(n)\eta(n)$ at small values of $n$, which means that positive (competitive) interactions dominate the average correlation. On the contrary we find that for large values of the abundances one has $P(n,n)>\eta(n)\eta(n)$: in this regime negative (mutualistic) interactions are the dominant ones. From the study of the connected probability $P(n,n)-\eta(n)\eta(n)$ it is therefore clear that strong non-Gaussianity affects (in a non trivial way) also the correlations when the disorder parameter $\hat{\sigma}$ is increased. We have then studied the function $P(n,m)-\eta(n)\eta(m)$ in the specific cases of either exclusively positive ($\alpha_{ij}>0$) or exclusively negative ($\alpha_{ij}<0$) interactions. Namely we have specialized the average of $P^{\alpha}(n,m)$ with respect to the sign of the interaction: